# #1#1#1# 1 (1)

### v1

@@ 3.2 Scalable Meta-Learning with Truncated Gradient Unrolling
4 Experiments
In this section, we empirically evaluate PERK’s long-context reasoning capabilities in two challenging scenarios: reasoning over (1) Needles-in-a-Haystack (NIAH, [32]) and (2) Drops-in-an-Ocean (DIO), a novel long context evaluation setting we propose.
Baselines   We compare our method to two long context reasoning adaptations applied to multiple models. In particular, we finetune transformer [60] and Mamba [21] models for prompting-based in-context reasoning (FT-ICR), as in conventional long-context finetuning [18, 19]. We apply FT-ICR to GPT-2 (127M parameters) [51], a small model with a limited context window (1024 tokens) and no long-context pretraining, Qwen-2.5-0.5B [50] (with a 32k token context window), and to two Mamba [21] models with sizes of 130M and 1.4B parameters. We refer to these baselines as FT-ICR (GPT-2), FT-ICR (Qwen), FT-ICR (Mamba-130M), and FT-ICR (Mamba-1.4B). Whenever evaluating tasks with contexts longer than the 1024 token limit of GPT-2, in order to effectively fine-tune it, we first provide GPT-2 with a series of additional positional embeddings instantiated using the interpolation method proposed in [29]. The Qwen models, meanwhile, all have sufficient context length for our experiments, and the Mamba models are designed to work with sequences of any length. In our experiments, we refer to models trained with PERK as PERK (GPT-2) and PERK (Qwen) depending on their base model. With only the LoRA parameters being trainable, PERK (GPT-2) contains 38M trainable parameters, and PERK (Qwen) has 140M trainable parameters.
4.1 Reasoning Over Needles-in-a-Haystack
Figure 2: Performance on NIAH with BabiLong and DIO with Student Records. All models are trained and tested on sequences ranging from 1K to 8K tokens. All PERK models are trained to first generate the supporting facts relevant to a query, followed by the final answer prediction. In contrast, baseline models directly generate the answer, as this approach yields better performance for them in this setting. The number of trainable parameters for each method is indicated in the legend.
Setup
We first evaluate PERK on reasoning over Needles-in-a-Haystack using the BabiLong framework [32] for assessing models’ reasoning abilities across relevant and distractor facts scattered in very long documents. We mainly focus on tasks involving single-hop (QA1), two-hop (QA2), and three-hop (QA3) reasoning. We use the framework to construct the tasks with long natural documents sampled from FineWeb-Edu [41], ensuring the content postdates our base model’s knowledge cutoff.
Results
In the top row of Figure 2, we show the evaluation results on BabiLong. PERK outperforms all baselines in all cases we test (across all models), including when compared to larger models. In particular, PERK (GPT-2) outperforms even FT-ICR (Qwen) and FT-ICR (Mamba-1.4B) for all task complexities and context lengths, with an average 20% performance gain. PERK (Qwen) achieves the highest accuracy across all task complexities (QA1, QA2, QA3) and context lengths (1K to 8K). PERK is also more robust to increasing context length and task complexity than the baselines. At best, PERK (Qwen) loses almost no accuracy with growing context length (for QA1 and QA2), and at worst (for QA3), it loses ${\sim}10\%$. When averaging across context lengths, PERK (Qwen)’s performance gain over FT-ICR (Qwen) increases from 10% to 41% when the number of reasoning hops increases from 1-hop (QA1) to 3-hop (QA3).
4.2 Reasoning Over Drops-in-the-Ocean
Needle-in-a-Haystack (NIAH) datasets for long-context evaluation typically have a fundamental limitation: the target information often exhibits stylistic differences from surrounding irrelevant text, creating an artificially simple test for identifying relevant information. In practical scenarios, relevant and irrelevant information will likely be distributionally similar, increasing the difficulty of identifying critical facts. As a result, we propose Drops-in-the-Ocean (DIO), a new evaluation setting that forms long contexts from structurally similar documents.
Setup
We construct a synthetic dataset, Student Records, where each context simulates a database containing multiple student records. Each record includes several attributes: ID, name, school, major, and grade. Unique IDs and names are generated for each database instance. Schools and majors are sampled with replacement from predefined pools, and grades are integers sampled uniformly from the range [0, 100]. Context length scales directly with the number of student records included. Inspired by BabiLong, we define evaluation tasks of varying complexity: Recall (retrieving attributes for a specific ID), Relation (comparing attributes between two IDs), and Aggregate (calculating the maximum, minimum, and average grade across all students). Crucially, to test generalization, specific entities (e.g., IDs, names) and attribute value combinations encountered in the test set are disjoint from those used during training. Examples are available in the Appendix.
Results
Results on the Student Records in Figure 2 mirror the trends seen in BabiLong. FT-ICR (GPT-2) fails to reason in the DIO setting with 0% accuracy or random performance (for Relation, the random performance is 50%). FT-ICR (Qwen) appears to be the strongest baseline, especially on recalling with a 1K context, where it briefly matches PERK (Qwen)’s near-perfect accuracy. PERK models again outperform all baselines. PERK (Qwen) and PERK (GPT-2) consistently achieve high accuracy (often >85%) and are more robust to increasing context length, even on the more challenging aggregation task with the 8k length, which requires the model to aggregate information across all 8k tokens. With Qwen as a base model, PERK achieves a 9% average gain over FT-ICR.
5 Analysis
5.1 Test-Time Length Generalization
We evaluate PERK’s ability to generalize to longer inputs than those seen in training [48].
Setup
We evaluate using the BabiLong tasks, which allow us to generate contexts of arbitrarily long lengths, and train a Qwen-2.5-0.5B base model using PERK on fixed-length contexts ranging from 1K to 8K tokens (as well as an FT-ICR baseline). Subsequently, we evaluate each model on contexts of up to 32K tokens, measuring each approach’s ability to generalize to new, longer lengths (i.e., extrapolation), as well as shorter contexts (i.e., interpolation). Overall, we train on contexts of increasing lengths (1K to 8K tokens) and test on contexts ranging from 1K to 32K tokens.
PERK generalizes to new context lengths at test time.
Figure 5 shows that PERK outperforms FT-ICR on the BabiLong QA1 (single-hop) and QA2 (two-hop) tasks for all generalization settings. For extrapolation, FT-ICR’s performance drops drastically as the inference context length exceeds the training context length, especially for shorter training contexts.
Figure 3: Test-time length extrapolation beyond 32K on BabiLong QA1 (a) and QA2 (b). Both PERK and FT-ICR are trained on 8K-token sequences. The context length for inference grows from 64K to 128K. PERK extrapolates substantially better than FT-ICR.
While PERK’s accuracy also decreases (from a higher starting point), the degradation is dampened at longer test lengths. For QA1 (Figure 5a), PERK trained on 1K sequences shows only a 42% drop on 32K sequences while FT-ICR exhibits a 52% decline. When trained on 8K sequences, PERK drops only 5% on 32K sequences, while FT-ICR still exhibits a 32% drop. Similar dynamics are observed on the more challenging QA2 task (Figure 5b). To stress-test the limit of PERK models’ length extrapolation ability, we test on sequences with context lengths beyond the context window of the Qwen-2.5 model – 32K tokens in Figure 3. Specifically, we test on sequences with tokens increasing from 64K to 128K, and demonstrate that PERK consistently extrapolates to longer contexts than the FT-ICR baseline. We acknowledge that with 128K token context windows, PERK’s accuracy experiences a noticeable drop to 61.4% for QA1 accuracy and 44.4% for QA2 accuracy. However, these scores remain substantially better than the 0% QA1 and QA2 absolute performance of the FT-ICR baselines at this same context length.
Meanwhile, for interpolation, FT-ICR performance drops on shorter contexts when finetuned on longer ones, matching observations from prior work [19]. In contrast, PERK maintains its performance and even increases performance for the QA2 task. Overall, PERK demonstrates strong length generalization, extrapolating to longer contexts while preserving high performance on shorter ones.
Figure 4: Test-time context length generalization evaluation on BabiLong QA1 (a) and QA2 (b): comparison between PERK and FT-ICR on the Qwen-2.5-0.5B model. The y-axis represents the training context lengths, while the x-axis indicates various test-time context lengths. We test for both interpolation (test lengths shorter than the training length) and extrapolation (test lengths longer than the training length). Bordered cells denote the boundary: evaluation on context lengths equal to those in training. PERK shows stronger generalization across both settings.
Figure 5: Positional Bias. Comparison of PERK and FT-ICR on 4K and 8K contexts, on Qwen-2.5-0.5B. We train on problems where the relevant information appears in the beginning (Pre), middle (Mid), or end (Post) of the context, and evaluate on all three positional settings. We also train models on contexts where the relevant information is randomly located (Rnd), testing these on all four positional distributions (Pre, Post, Mid, Rnd). Bordered cells show in-distribution performances. PERK demonstrates strong positional robustness.
5.2 Robustness to Positional Biases
Prior work has shown that the position of relevant information in long contexts affects a language model’s ability to utilize that information [39]. In this experiment, we test PERK’s robustness to positional biases by evaluating its test-time performance on contexts where the position of relevant information is distributed differently from those seen in training.
Setup
We use API documents from [45], where each document has an associated API call, and create long contexts containing multiple API-call-to-document pairs. In this task, the model must retrieve the correct API call based on user instructions and documents. Using fixed 4k and 8k token contexts, we train PERK and FT-ICR models on contexts with the target document at varying positions and test them on contexts distributed across all positions.
PERK generalizes regardless of information position in the context.
Figure 5 compares PERK to FT-ICR under different position distributions (at train and test time) and context lengths. When trained on contexts with relevant documents randomly located (Rnd) throughout the context, PERK outperforms FT-ICR by 24%. Furthermore, FT-ICR shows large performance drops when the relevant document consistently appears at different locations (Pre, Mid, and Post) at test time. Meanwhile, shifting relevant positions at test-time shows minimal effect (within 1-2%) on PERK.
To further stress-test positional generalization, we also force relevant documents into specific positions during training and testing, spanning the full context. We find that FT-ICR easily overfits to the position pattern, completely failing to generalize to test-time position changes (performance drops to close to 0% when the position shifts at test time).
Although PERK underperforms FT-ICR when the train-test positions are the same and the beginning (Pre) or end (Post) of the context, performance remains consistent regardless of the positional distribution shift between train and test time. We attribute PERK’s robustness in this setting to the fact that documents in the long context are encoded as members of a batch into the parameter space, limiting the overall impact of absolute position in the sequence. Meanwhile, FT-ICR directly aligns specific positions with answers during training.
Figure 6: (a): Peak GPU memory usage during training with context lengths ranging from 1K to 8K tokens for Reckoning and PERK. While PERK successfully scales to 8K tokens, Reckoning encounters out-of-memory (OOM) errors at shorter context lengths. (b): Memory footprint and wall-clock runtime during inference as context length increases (up to 128K tokens), comparing PERK with FT-ICR. Curves that terminate before 128K indicate that the method failed with an OOM error at longer context lengths, preventing further measurement. PERK demonstrates more efficient scaling in both memory and runtime, particularly for extremely long sequences.
5.3 Training and Inference Complexity Analysis
We first show PERK’s training scalability compared to prior test-time learning method Reckoning [13] by measuring each method’s peak training memory across different context lengths. We then evaluate PERK’s training efficiency on long contexts by measuring its inference memory cost and run-time, compared to in-context reasoning with finetuned models.
Setup
We complete all runs on a single Nvidia H100 GPU with 93 GB of VRAM. We perform four gradient update steps for the test-time and inner loop adaptations. For training, we have four truncation levels: truncate 0 (basically gradient unrolling without truncation), 1, 2, and 3 steps. For inference, both methods generate 64 tokens using greedy decoding with Huggingface’s text generation pipeline. We measure runtime complexity in seconds (wall-clock time) and memory complexity in gigabytes. Both methods undergo 10 hardware warm-up iterations.
As explained in Section 3, PERK splits long sequences into batches of shorter sequences. Here, each batch has an effective context length defined as 128 tokens. Since PERK performs gradient-based encoding over these batches, we can apply gradient accumulation to trade runtime for reduced memory consumption. We thus analyze PERK across varying gradient accumulation steps, ranging from 2 to 16. Note that for the minimum context length of 1024 tokens, which yields a batch of 8 sequences, gradient accumulation with 16 steps is infeasible as it exceeds the batch size.
Figure 6a shows that, as the training context length grows from 1K to 8K tokens, Reckoning quickly runs into Out-Of-Memory (OOM) errors at 2K, while PERK with increasing truncation steps continues to fit the maximum GPU memory. We successfully validate PERK’s stronger scalability on long contexts.
Figure 6b for inference complexity shows that PERK provides more efficient memory and runtime scaling for extremely long contexts compared to FT-ICR. While FT-ICR is initially more efficient, its memory and runtime grow rapidly, leading to OOM errors at a context length of 128K. In contrast, PERK can manage the long sequences through gradient accumulation, which, while increasing runtime, reduces the memory footprint. For instance, at 8K tokens, increasing the accumulation steps from 1 to 16 reduces memory usage from 35.2GB to 5.9GB, although the runtime increases from 1.9s to 8.5s. Crucially, even with this runtime trade-off, PERK runs faster than FT-ICR at very long contexts where both can operate. At 64K tokens, FT-ICR takes 32.6s and 55.7GB, whereas PERK with 16 steps of accumulation completes in 11.4s, using only 19.6GB. With 8 steps, it is even faster at 9.7s, using 35.2 GB. Ultimately, at 128K tokens, where FT-ICR fails, PERK with 16 steps successfully processes the context using 35.2GB in 20.9s, showing that PERK provides a practical path to handle extreme context lengths efficiently in both memory and runtime when compared to standard approaches.
6 Related Work
Test-Time Learning
MAML [17] introduces optimization-based model-agnostic meta-learning for few-shot adaptation. [13] applies MAML to GPT models, integrating knowledge into model parameters through test-time gradient updates. Some methods achieve test-time learning in the model architecture. [58] proposes RNN layers with learnable hidden states updated via self-supervised gradients on test inputs. Titans [8] incorporates differentiable memory with transformers for rapid inference-time memorization. Building on the Titan memory, ATLAS [7] proposes a new transformer architecture that contains the differentiable memory module as part of the network to support large-scale distributed pretraining. Other approaches adapt through context alone. MetaICL [43] meta-learns from in-context examples without gradient updates. Our method, PERK, introduces parameter-efficient meta-learning using lightweight adapters, improving scalability for long contexts.
Long-Context Language Models
Developing long-context models has become a major research topic of its own. Many works explore extending transformer-based language models’ context windows with minimal training, either by position extrapolation [12, 47, 63, 69] or improving the transformer’s attention mechanism [9, 44, 56, 27]. Other works show that using the default attention, applying simple position extrapolation, and fine-tuning the model on long documents [18, 42] or synthetically-generated long data [1, 64] yields much stronger results. The FT-ICR baseline in our work follows this paradigm. Other LLM works [37, 20, 19] propose achieving long-context capabilities through a long-context continued training stage between standard pre-training and supervised fine-tuning. In contrast to all the approaches above, PERK achieves long-context reasoning through the ability to encode long sequences using test-time gradient updates on a lightweight model adapter.
Long-Context Architectures
In parallel, there have been many efforts in designing more efficient architectures, for example, linear attention with RNNs [11, 46, 21, 15, 66, 6] and alternative attention architectures [53, 59, 22]. However, these new architectures often require training from scratch, and many have inherent limitations in terms of long-context recall [26, 3]. Recent works also explore hybrid models [37, 16] or distilling existing LLMs into hybrid models [61] and show promising results. In comparison, PERK directly augments off-the-shelf pre-trained LLMs without requiring additional architecture modifications or extensive pre-training, while naturally being parameter-efficient.
Long-Context Evaluation
Many benchmarks have been proposed to evaluate language models’ long-context abilities [31, 55, 4, 71, 23, 70, 68, 62]. Some works focus on the impact of extending input lengths with irrelevant contexts [32, 35] or changing relevant information positions [39] on LLM reasoning. Many works also reframe natural-language-processing tasks, such as retrieval [34, 49], summarization [30], and in-context learning [36, 65, 10], as long-context challenges. In this work, we define Drops-in-the-Ocean as a new long-context evaluation paradigm that makes the long context as distributionally similar to relevant information as possible.
7 Conclusion
In this work, we introduce PERK, a parameter-efficient meta-learning approach that learns to perform long-context reasoning by encoding information into a lightweight adapter through test-time gradient updates. Our experiments on multiple long-context benchmarks (Needle-in-the-Haystack, Drops-in-the-Ocean, and API retrieval) demonstrate substantial performance gains over models finetuned for classical in-context reasoning across reasoning complexity and context lengths. Our generalization analysis highlights PERK’s strong test-time length extrapolation ability, where the model can generalize to unseen contexts 32 times longer than the training data. When relevant information shifts positions at test-time, PERK robustly generalizes across different positions. Finally, we show that PERK scales efficiently at inference to extremely long contexts. Overall, PERK excels in performance, generalization, and efficiency for long-context reasoning.
@@ A.1 BabiLong
We construct the train and evaluation data using the data generation framework proposed by [32]. As we mentioned in the experiment setup, we mainly focus on tasks involving single-hop (QA1), two-hop (QA2), and three-hop (QA3) reasoning.
@@
To build the long context, we use the framework to construct the tasks with long natural documents sampled from FineWeb-edu [41], specifically from the CC-MAIN-2024-42 subset, ensuring the content postdates our base model’s knowledge cutoff. Given a particular context window, such as 8192, we sample that many tokens from the corpus as the Haystack long context.
@@ A.2 Student Records
Aggregation is the most difficult task in this dataset. This task requires many-hop reasoning ability, where the model needs to aggregate attributes from all students in a long context to make an accurate conclusion. We focus on three types of numerical aggregation among the students’ grades: (1) Maximization, (2) Minimization, and (3) Averaging. For example, we ask the model “What is the average grade of all students”, the model should be able to compute the sum of all student grades, count the number of total students in the context, and then compute the average, “49.3”.
Figure 7: Dataset Examples. Here we show an example from the three datasets we used: (Top) BabiLong, (Middle) Student Records, and (Bottom) Lost-in-the-Middle API Retrieval.
@@
A.3 Lost-in-the-Middle API Retrieval
To analyze a model’s robustness against relevant information positions, we follow the setting proposed by Lost in the Middle [39]. We use API documents from [45], where each document has an associated API call, and create long contexts containing multiple API-call-to-document pairs. We place the relevant API document in a particular user instruction at different locations in the context (beginning, middle, end, and random position).
@@ Appendix B Training and Inference Details
Each training (for PERK and baselines) sets 2 epochs as the maximum runtime with early stopping based on validation loss. In practice, most training runs typically stop early, within one epoch of runtime. We set the baseline training and PERK’s outer loop learning rate to 1e-5, and the weight decay to 0.01. We use a cosine learning rate scheduler with 0.03 of the total optimization steps as the warm-up phase. We use PyTorch Lightning  [^1] https://lightning.ai/docs/pytorch/stable/ as the trainer for PERK. The Qwen and GPT-2 FT-ICR baselines are trained using the Open-Instruct [33] trainer from AI2, and the Mamba FT-ICR baselines are trained using the Huggingface Trainer  [^2] https://huggingface.co/docs/transformers/en/main_classes/trainer. For the stress test of length generalization beyond 64K tokens, we change the default maximum position of 32768 to 131072.
@@
We set LoRA rank to 256 for both PERK GPT-2 and PERK Qwen. We use the default values for the other hyperparameters, including setting the alpha to 16 and the dropout to 0.1. We apply LoRA to all modules in the model architecture. For GPT-2, the number of active training parameters of PERK is 38M, 30% of the total parameter number. For Qwen, the number of active training parameters of PERK is 140M, which is also 30% of the total parameter number. We set the scaling factor $\alpha$ to 16 initially and enable rank-stabilized LoRA fine-tuning [28], which results in a target $\alpha$ of 256, corresponding to a rank of 256.
@@
For the inner loop adaptation, we use 4 inner loop gradient steps. For the truncated gradient unrolling, we truncate steps 1 and 2 for input context length $\leq$4K, and we truncate steps 1, 2, and 3 for context length $>$ 8K. We follow [67]’s practice on how to train a good MAML model and use per-layer-per-step dynamic learning rate for the inner loop, using 5e-5 as the initial value. For the language modeling loss in the inner loop, we use a weighted version similar to CaMeLS [25]. The token-level weights are predicted via a weighting model implemented as a two-layer MLP network. The parameters of the weighting model are co-meta-learned along with the LoRA parameters and the adaptive per-layer-per-step learning rates. We use the differentiable version of the AdamW [40] optimizer for the inner loop optimization, implemented by the Higher  [^3] https://github.com/facebookresearch/higher library for higher-order optimization. We compute higher-order derivatives using the automatic differentiation mechanism implemented by PyTorch [^4] https://docs.pytorch.org/tutorials/beginner/blitz/autograd_tutorial.html. The outer loop optimizer is the standard PyTorch implementation of AdamW.
@@
At inference time, we generate the answer for a reasoning problem via greedy decoding (temperature is set to 0). The maximum number of tokens that can be generated is 512. The EOS token is set to the default one defined in the tokenizer. The generation is handled by the text-generation pipeline implemented by Huggingface  [^5] https://huggingface.co/docs/transformers/v4.43.4/en/main_classes/pipelines.

### v2

@@ 7 Conclusion
Acknowledgments
We gratefully acknowledge the support of the Swiss National Science Foundation (No. 215390), Innosuisse (PFFS-21-29), the EPFL Center for Imaging, Sony Group Corporation, and a Meta LLM Evaluation Research Grant.
@@ A.1 BabiLong
We construct the train and evaluation data using the data generation framework proposed by kuratov2024babilong. As we mentioned in the experiment setup, we mainly focus on tasks involving single-hop (QA1), two-hop (QA2), and three-hop (QA3) reasoning.
@@
To build the long context, we use the framework to construct the tasks with long natural documents sampled from FineWeb-edu (lozhkov2024fineweb-edu), specifically from the CC-MAIN-2024-42 subset, ensuring the content postdates our base model’s knowledge cutoff. Given a particular context window, such as 8192, we sample that many tokens from the corpus as the Haystack long context.
@@ A.2 Student Records
Aggregation is the most difficult task in this dataset. This task requires many-hop reasoning ability, where the model needs to aggregate attributes from all students in a long context to make an accurate conclusion. We focus on three types of numerical aggregation among the students’ grades: (1) Maximization, (2) Minimization, and (3) Averaging. For example, we ask the model “What is the average grade of all students”, the model should be able to compute the sum of all student grades, count the total number of students in the context, and then compute the average, “49.3”.
Figure 7: Dataset Examples: Here we show an example from the three datasets we used: (Top) BabiLong, (Middle) Student Records, and (Bottom) Lost-in-the-Middle API Retrieval. PERK first encode the contexts into the memory scratchpad through self-supervised test-time learning. Next, the user can ask any number of questions about that context, and the LLM can reason over the updated memory scratchpad to generate the answers.
@@
A.3 Open-Domain Multi-doc QA
We select HotpotQA (yang-etal-2018-hotpotqa) and TriviaQA (joshi-etal-2017-triviaqa) as the main evaluation for open-domain QA over long-context documents. HoptpotQA tests the model’s ability to perform 2-hop reasoning over two gold documents randomly located in a long list of documents. TriviaQA tests the model’s ability to do reading comprehension over very long documents where the target information can be scattered around the long context. We refer to the original papers for QA construction details, data sources, and task examples of these two benchmarks.
Build Long Context
To build long contexts for these two benchmarks, given a specific context window (for example, 32768 tokens), we follow yen2025helmetevaluatelongcontextlanguage’s process for their RAG evaluation. We extend the original context by injecting retrieved distractor documents (zhang2024mgte) at random positions until reaching the context window limit.
A.4 Lost-in-the-Middle API Retrieval
To analyze a model’s robustness against relevant information positions, we follow the setting proposed by Lost in the Middle (liu2024lost). We use API documents from patil2024gorilla, where each document has an associated API call, and create long contexts containing multiple API-call-to-document pairs. We place the relevant API document in a particular user instruction at different locations in the context (beginning, middle, end, and random position).
@@ Appendix B Training and Inference Details
Each training (for PERK and baselines) sets 2 epochs as the maximum runtime with early stopping based on validation loss. In practice, most training runs typically stop early, within one epoch of runtime. We set the baseline training and PERK’s outer loop learning rate to 1e-5, and the weight decay to 0.01. We use a cosine learning rate scheduler with 0.03 of the total optimization steps as the warm-up phase. We use PyTorch Lightning  [^1] https://lightning.ai/docs/pytorch/stable/ as the trainer for PERK. The Qwen and GPT-2 FT-ICR baselines are trained using the Open-Instruct (lambert2025tulu3pushingfrontiers) trainer from AI2, and the Mamba FT-ICR baselines are trained using the Huggingface Trainer  [^2] https://huggingface.co/docs/transformers/en/main_classes/trainer. For the stress test of length generalization beyond 64K tokens, we change the default maximum position of 32768 to 131072.
@@
We set LoRA rank to 256 for both PERK GPT-2 and PERK Qwen. We use the default values for the other hyperparameters, including setting the alpha to 16 and the dropout to 0.1. We apply LoRA to all modules in the model architecture. For 1B scale and below, we apply LoRA to all layers in the model. However, for 7B and 8B models, we apply it only to the top 4 layers in order to reduce the memory overhead of meta-learning. We set the scaling factor $\alpha$ to 16 initially and enable rank-stabilized LoRA fine-tuning kalajdzievski2023rslora, which results in a target $\alpha$ of 256, corresponding to a rank of 256.
@@
For the inner loop adaptation, we use 4 inner loop gradient steps. For the truncated gradient unrolling, we truncate steps 1 and 2 for input context length $\leq$4K, and we truncate steps 1, 2, and 3 for context length >> 8K. We follow ye2022how’s practice on how to train a good MAML model and meta-learn a set of per-layer-per-step dynamic learning rates for the inner loop as part of the outer loop learning, using 5e-5 as the initial value. This approach learns a separate learning rate for each network layer and each adaptation step in the inner loop. These learning rates are optimized during the meta-training process alongside the model parameters. Note that these per-layer-per-step learning rates are learned parameters, not hyperparameters that require manual tuning. During meta-training, our algorithm automatically learns to adjust these learning rates for each layer and inner loop step dynamically based on the meta-objective. This approach reduces the hyperparameter tuning burden compared to manually setting the adaptive learning rates. For the language modeling loss in the inner loop, we use a weighted version similar to CaMeLS (hu2023onlineadaptation). The token-level weights are predicted via a weighting model implemented as a two-layer MLP network. The parameters of the weighting model are co-meta-learned along with the LoRA parameters and the adaptive per-layer-per-step learning rates. We use the differentiable version of the AdamW (loshchilov2017decoupled) optimizer for the inner loop optimization, implemented by the Higher  [^3] https://github.com/facebookresearch/higher library for higher-order optimization. We compute higher-order derivatives using the automatic differentiation mechanism implemented by PyTorch [^4] https://docs.pytorch.org/tutorials/beginner/blitz/autograd_tutorial.html. The outer loop optimizer is the standard PyTorch implementation of AdamW.
Long-context Data Collation for PERK
At training time, to collate a reasoning problem with a long context and a set of associated question-answer pairs, we first convert a long context into a batch of subsequences, with relatively even lengths. In practice, we set the effective context length (i.e., the number of tokens of a subsequence on average) as 256, and divide the long context into $N$ chunks where each chunk is approximately 256 tokens long. For a long context with 8192 tokens, the batch size is thus $N=32$. PERK’s inner loop adaptation encodes the batch of subsequences in parallel through gradient updates. At inference time, if we want to test PERK on a long context with 32K tokens, the effective context length stays the same as approximately 256 tokens, while the batch size increases to $N=128$.
Note that splitting a long context into a batch of shorter sequences can either be permutation-invariant or maintain the original ordering. PERK induces ordering through explicit indexing in the prompt formatting, which allows the model to learn implicit ordering during the encoding phase. Specifically, when chunking a long context, we can preserve ordering information by indexing each chunk. For example, if we split a long document into segments A, B, C, D, E, we format them as: ["Document 1: A", "Document 2: B", …, "Document 5: E"]. During the inner loop adaptation with the NLL objective, the LoRA parameters learn to associate these indices with their content, effectively encoding the ordering information.
@@
At inference time, we generate the answer for a reasoning problem via greedy decoding (temperature is set to 0). The maximum number of tokens that can be generated is 512. The EOS token is set to the default one defined in the tokenizer. The generation is handled by the text-generation pipeline implemented by Huggingface  [^5] https://huggingface.co/docs/transformers/v4.43.4/en/main_classes/pipelines. For the length generalization analysis, we apply recent training-free methods Yarn (peng2023yarnefficient) and Dual Chunk Attention (DCA, an2024dca) to the finetuned Qwen models. We use vLLM  [^6] https://docs.vllm.ai/en/latest (0.10.2) as the inference engine to apply Yarn + DCA. For Yarn, we apply a factor of 4.0 to the original max position embeddings, 32768, and scale the context window to 131072. For DCA, we use vLLM’s DUAL_CHUNK_FLASH_ATTN backend for generation.